I hate these.
Mr. mcgoo is 8times as old as his son, Gomer. In 3 years he will be 5 times his sones age. How old is gomer now?
Thank you kind person

Answer by Earlsdon(6294)

Let M = McGoo's present age and G = Gomer's present age.
From the problem, you can write:
1) M = 8G "McGoo is 8 times as old as his son, Gomer"
2) M+3 = 5(G+3) "In three years, he (M) will be five times his son's (G's) age"
Substitute the M = 8G from equation 1) into equation 2) and solve for G.
2a) (8G)+3 = 5(G+3) Simplify.
8G+3 = 5G+15 Subtract 5G from both sides.
3G+3 = 15 Subtract 3 from both sides.
3G = 12 Finally, divide both sides by 3.
G = 4
Gomer is 4 years old.